(a) A 1.06 kg mass is at the center of a uniform spherical shell of mass 1.10 1020 kg and radius 1.01 106 m. Find the resultant gravitational force on the mass.

(b) Find the resultant gravitational force on the mass if it is placed just outside the shell. magnitude =

(c) Find the value of g just outside the shell.
Answer #1

a) Fnet = 0

why  

because this small mass is situated at the center and we know that gravitational force is always attractive

so one half of the shell and other halves of the shell apply the same amount of force but in opposite direction, so net force is zero

(b))

just out side

F = GMm/r2

= 6.67*10^-11 *1.06*1.10*10^20 /(1.01*10^6)^2 =

1.1 x1020 6.67x 1011.06: (1.01x 106)2

= 7.62397804e-3 N answer

(c)

g = GM/r2

= 6.67*10^-11 *1.10*10^20 /(1.01*10^6)^2

  1.1 1020 6.67x 10-11 (1.01 1062

g = 7.19243211e-3 m/s2

****************************************************************************
Goodluck Comment in case any doubt, will reply for sure..